# for Dynamic Neural Radiance Fields

### v2

@@ IV-C Novel View Synthesis Quality
On the BONN RGB-D dataset, COLMAP is challenged on several sequences (cf. Tab. IV). DynaMoN can retrieve camera poses for all scenes (Tab. II) and performs well on novel view synthesis for all sequences (Tab. IV). In addition, the PSNR is increased by the alternating training scheme. However, for the cases where COLMAP succeeds in the pose estimation, the camera poses from SfM also lead to decent novel view synthesis results. Examples of renderings can be seen in Fig. 4.
IV-D Robustness on Sparse Trajectories
Finally, we evaluate the robustness of different parts of our pose retrieval when faced with a dynamic and sparse trajectory. Using only every 10th frame of the crowd2 sequence, we show how the DROID-SLAM poses compare to the ones from the dynamics-masked camera pose retrieval and our neural camera pose refinement in terms of novel view rendering quality. Looking at the PSNR values in Table V, it can be noted that each of our proposed components incrementally increases the reconstruction quality. Thus, when confronted with a sparse trajectory of a highly dynamic scene, DynaMoN outperforms pure DROID-SLAM making it more robust to realistic, real-world scenarios with imperfect camera recordings.
@@ V Discussion
In comparison to traditional sparse reconstructions, our approach addresses camera localization and novel view synthesis in combination. While this leads to more pleasing 3D visualizations and novel view synthesis results, the computing time is higher compared to traditional camera localization and 3D representation methods.
For camera pose estimation, our approach shows an improved performance compared to the state-of-the-art. Only using our motion masks already provides an improved performance. Moreover, for scenes with large motions, the combination with semantic segmentation masks provides significantly improved camera localization results. The results presented in Table III indicate that RoDynRF [33], despite being a state-of-the-art method that integrates combined camera regression with novel view synthesis, faces significant challenges in handling highly dynamic camera movements and dynamic scenes. In contrast, our method demonstrates superior robustness and yields improved outcomes when compared to both RoDynRF [33] and HexPlane [4] employing COLMAP poses on the demanding TUM RGB-D dataset. Furthermore, our iterative pose refinement process enhances robustness, as evidenced in Table V, and achieves better performance in novel view synthesis on both datasets.

### v3

@@ IV-D Novel View Synthesis Quality
On the BONN RGB-D Dynamic dataset, COLMAP is challenged on several sequences (cf. Tab. IV). DynaMoN can retrieve camera poses for all scenes (Tab. II) and performs well on novel view synthesis for all sequences (Tab. IV). In addition, the PSNR is increased by the alternating training scheme. However, for the cases where COLMAP succeeds in the pose estimation, the camera poses from SfM also lead to decent novel view synthesis results. Examples of renderings can be seen in Fig. 4.
IV-E Robustness on Sparse Trajectories
In this section, we evaluate the robustness of different parts of our pose retrieval when faced with a dynamic and sparse trajectory. Using only every tenth frame of the crowd2 sequence, we show how the DROID-SLAM poses compare to the ones from the dynamics-masked camera pose retrieval and our neural camera pose refinement in terms of novel view rendering quality. Looking at the PSNR values in Table V, it can be noted that each of our proposed components incrementally increases the reconstruction quality. Thus, when confronted with a sparse trajectory of a highly dynamic scene, DynaMoN outperforms pure DROID-SLAM making it more robust to realistic, real-world scenarios with imperfect camera recordings.
IV-F Runtime and Memory Evaluation
Finally, the runtime of DynaMoN is measured and the memory requirements are mentioned. Both are presented in comparison with RoDynRF [34] on a workstation with an NVIDIA RTX 4090 graphics card with 24 gigabytes of video random-access memory (VRAM). We report the results for the TUM RGB-D fr2/desk-person sequence in Table VI as it is by far the longest scene of both datasets comprising 4 067 frames. In terms of memory requirement, RoDynRF can only be trained with a batch size of 512 in this setting, while DynaMoN is more memory-efficient, enabling a batch size of 1 024. From the experiments, it can be concluded that the total time for pre-processing, training and rendering of DynaMoN is comparable to RoDynRF if the default training strategy with 200 000 iterations is utilized. However, if our model is trained for 100 000 iterations which is the default value of RoDynRF, the total time can be greatly reduced without a significant drop in photometric accuracy. For this shorter training, the NeRF and pose optimization of DynaMoN alternate every 2 500 iterations. Furthermore, a decrease of the batch size to the same value as in the competitor method further speeds up the training time at the cost of a longer rendering time while staying comparably accurate. It has to be noted that the pre-processing of RoDynRF consisting of monocular depth prediction, optical flow prediction, motion mask generation and dataloading takes about eight times longer than our pre-processing loading the data and generating initial poses as well as motion masks.
@@ V DISCUSSION
In comparison to traditional sparse reconstructions [15, 14, 16], our approach addresses camera localization and dynamic novel view synthesis in combination. While this leads to more pleasing 3D visualizations and novel view synthesis results, the computing time is higher compared to traditional camera localization and non-dynamic 3D representation methods [9, 10, 11]. Nevertheless, compared to combined dynamic camera localization and dynamic novel view synthesis approaches [34], our approach requires only 60% of the training time when using the same number of iterations and batch size, while being more robust in more dynamic scenes with dynamic camera motion.
Moreover, our approach shows an improved performance compared to the state-of-the-art for camera localization. Only using our motion masks already improves the performance. Moreover, for scenes with large motions, the combination with semantic segmentation masks provides significantly improved camera localization results.
The results presented in Table III indicate that RoDynRF [34], despite being a state-of-the-art method that integrates combined camera regression with novel view synthesis, faces significant challenges in handling highly dynamic camera movements and dynamic scenes. In contrast, our method demonstrates superior robustness and yields improved outcomes when compared to both RoDynRF [34] and HexPlane [4] employing COLMAP poses on the demanding TUM RGB-D dataset. Furthermore, our iterative pose refinement process enhances robustness, as evidenced in Table V, and achieves better performance in novel view synthesis on both datasets.